lim

x→0

x

2

sin 


1
Expert's answer
2021-08-26T14:36:19-0400

"\lim\limits_{x\to0}x^2\sin(\dfrac{1}{x})"

"-1\leq\sin(\dfrac{1}{x})\leq1"

"-x^2\leq x^2\sin(\dfrac{1}{x})\leq x^2"

"\lim\limits_{x\to0}x^2=\lim\limits_{x\to0}(-x^2)=0"

Then by Squeeze Theorem


"\lim\limits_{x\to0}x^2\sin(\dfrac{1}{x})=0"